a new elementary algebra textbook sells for $85 and a used elementary algebra textbook sells for $65. How many new textbooks are sold if the receipts for 125 new and used textbooks is $9805?
Answer by fractalier(6550)

Call the numbers of new and old textbooks, x and y.
Then the setup looks like this...
x + y = 125 and the value equation is
85x + 65y = 9805
Now multiply the top one by 85 and subtract...we get
85x + 65y = 9805
-(85x + 85y = 10625)
---------------------
-20y = -820
Now divide by -20 and get
y = 41 old textbooks and so
x = 84 new textbooks